OpenStudy (anonymous):

Factor theorem Let P(x) be a polynomial Prove that a.)If P(c)= 0 then x-c is a factor of P(x) b.) If x-c is a factor of P(x then P(c)= 0

OpenStudy (anonymous):

@dumbcow pls help

OpenStudy (dumbcow):

hmm not sure how to about the formal proof...since P(c) = 0 , c is a zero thus (x-c) is factor of P(x)

OpenStudy (dumbcow):

you may want to try tagging someone else on this one...i plead to common sense but that doesn't work in mathematics :)

OpenStudy (debbieg):

Can you use the remainder theorem? If you divide a degree n polynomial P(x) by (x-c), you get: P(x)=(x-c)*Q(x)+r where the remainder r is a constant, and Q(x) is a polynomial of degree n-1. If you can use that, then (1) should follow pretty easily when you assume that P(c)=0 (just show that the remainder in the division by x-c=0, then that means that x-c is a factor).